An RL circuit with L = 5.00 H and an RC circuit with C = 6.00
Answer #1

series RL
time constant = L / R

series RC
time constant = R * C

L/R = R*C and R^2 = L / C = 5/ 6* 10^-6
R^2 = 0.833 * 10^6
R = 912.68 ohms

b)
time constant = R * C = 912.68 * 6x10^-6= 5476.13 10^-6 = 5.47613* 10^-3 s
time constant= 5.47613x10^-3 S = 5.47613 mS
Answer #2

L/R= R*C


so R^2=L/C


R=912.87 ohms


b)time constant=R*C=5.477 ms
Answer #3

Time constant t = CR for the RC circuit and t = L/R for the RL circuit


CR = L/R rearranges to


R^2 = L/C = 5.00/6.0*10^-6


R = 912 ohms


t = CR = 5.4* 10^-3 s
Answer #4

L/R=RC
R^2=L/C=0.83 MEGA ohm
a)R=0.913 kohm
b)Time constant=L/R=5/0.913 kohm=5.48 milli seconds